# Risk-Aware Adaptive Control Barrier Functions for Safe Control of Nonlinear Systems under Stochastic Uncertainty

### v1

@@ Abstract
This paper addresses the challenge of ensuring safety in stochastic control systems with high-relative-degree constraints while maintaining feasibility and mitigating conservatism in risk evaluation. Control Barrier Functions (CBFs) provide an effective framework for enforcing safety constraints in nonlinear systems. However, existing methods struggle with feasibility issues and multi-step uncertainties. To address these challenges, we introduce Risk-aware Adaptive CBFs (RACBFs), which integrate Discrete-time Auxiliary-Variable adaptive CBFs (DAVCBFs) with coherent risk measures. DAVCBFs introduce auxiliary variables to improve the feasibility of the optimal control problem, while RACBFs incorporate risk-aware formulations to balance safety and risk evaluation performance. By extending discrete-time high-order CBF constraints over multiple steps, RACBFs effectively handle multi-step uncertainties that propagate through the system dynamics. We demonstrate the effectiveness of our approach on a stochastic unicycle system, showing that RACBFs maintain safety and feasibility while reducing unnecessary conservatism compared to standard robust formulations of discrete-time CBF methods.
@@ II-A Discrete-Time High-Order CBFs
The output $y_{t}=h(\mathbf{x}_{t})$ of system (1) is said to have relative degree $m$ if
@@
i.e., $m$ is the number of steps (delay) in the output $\mathbf{y}_{t}$ in order for any component of the control input $\mathbf{u}_{t}$ to explicitly appear ($\textbf{0}_{q}$ is the zero vector of dimension $q$).
@@
where $\alpha(\cdot)$ is defined linear and $0<\gamma_{i}\leq 1,i\in\{1,\dots,m\}$.
@@ II-B Coherent Risk Measure
The one-step conditional risk measures $\rho_{t}:\mathcal{H}_{t+1}\to\mathcal{H}_{t}$, $t=1,\dots,N-1$, as in (3), are called a coherent risk measure if they satisfy the following properties:
@@
Conditional Value-at-Risk (CVaR) Let $h\in\mathcal{H}$ be a stochastic variable where higher values indicate safer performance. Since we aim to prevent $h$ from becoming negative (as seen in Eq. (4)), we treat negative values of $h$ as a loss and measure left-tail risk using Conditional Value-at-Risk (CVaR) or Value-at-Risk (VaR) at the lower quantile. For a confidence level $\beta\in(0,1)$, VaR is defined as:
@@ III Problem Formulation and Approach
A control policy is feasible if all constraints derived from previously mentioned requirements are satisfied and mutually non-conflicting for all $0\leq t\leq t_{T}$. In this paper, we consider the following problem:
@@
Find a feasible control policy for system (13) such that the previously mentioned requirements are satisfied and cost (14) is minimized.
@@ IV-A Enhancing Optimization Feasibility with DAVCBFs
$\epsilon$ is a positive constant which can be infinitely small, the auxiliary input $\nu_{i,t}$ will appear in $\varphi_{i,0}(\bm{\pi}_{i,t+j_{m}})$ and $j_{m}=m+1-i$.
@@
Equation (19) and $\varphi_{i,m+1-i}(\bm{\pi}_{i,t})\geq\epsilon$ are equivalent. If this inequality constraint is satisfied, we have $\bigtriangleup\varphi_{i,m-i}(\bm{\pi}_{i,t})+\alpha_{i,m+1-i}(\varphi_{i,m-i}(\bm{\pi}_{i,t}))\geq\epsilon$. Rewrite this, we have
@@
$\alpha_{j}(\psi_{j-1}(\mathbf{x}_{t},\bm{\Pi}_{t}))\leq\psi_{j-1}(\mathbf{x}_{t},\bm{\Pi}_{t})$, $\forall(\mathbf{x}_{t},\bm{\Pi}_{t})\in\mathcal{C}_{0}\cap,...,\cap\mathcal{C}_{m-1}$ and each $a_{i,t}>0,i\in\{1,\dots,m\}$. Note that the input $\mathbf{u}_{t}$ and auxiliary input $\bm{\nu}_{t}$ will appear in $\psi_{0}(\mathbf{x}_{t+m},\bm{\Pi}_{t+m})$ and Eq. (25) is equivalent to $\psi_{m}(\mathbf{x}_{t},\bm{\Pi}_{t})\geq 0$.
@@
Given a DAVCBF $h(\mathbf{x})$ from Def. 5 with corresponding sets $\mathcal{C}_{0},\dots,\mathcal{C}_{m-1}$ defined by (23), if $(\mathbf{x}_{0},\bm{\Pi}_{0})\in\mathcal{C}_{0}\cap\dots\cap\mathcal{C}_{m-1},$ then if there exists solution of Lipschitz controller $(\mathbf{u}_{t},\bm{\nu}_{t})$ that satisfies the constraint in (25) and also ensures $(\mathbf{x}_{t},\bm{\Pi}_{t})\in\mathcal{C}_{m-1}$ for all $t\in\mathbb{N}_{\geq 0}$, then $\mathcal{C}_{0}\cap\dots\cap\mathcal{C}_{m-1}$ will be rendered forward invariant for system (1), $i.e.,(\mathbf{x}_{t},\bm{\Pi}_{t})\in\mathcal{C}_{0}\cap\dots\cap\mathcal{C}_{m-1},\forall t\in\mathbb{N}_{\geq 0}$. Moreover, $h(\mathbf{x}_{t})\geq 0$ is ensured for all $t\in\mathbb{N}_{\geq 0}$
@@
If $h(\mathbf{x})$ is a DAVCBF, then satisfying constraint in (25) while ensuring $(\mathbf{x}_{t},\bm{\Pi}_{t})\in\mathcal{C}_{m-1}$ for all $t\in\mathbb{N}_{\geq 0}$ is equivalent to make $\psi_{m-1}(\mathbf{x}_{t},\bm{\Pi}_{t})\geq 0,\forall t\in\mathbb{N}_{\geq 0}$. Since $a_{m,t}>0$, we have $\frac{\psi_{m-1}(\mathbf{x}_{t},\bm{\Pi}_{t})}{a_{m,t}}\geq 0.$ Based on (22), since $(\mathbf{x}_{0},\bm{\Pi}_{0})\in\mathcal{C}_{m-2}$ (i.e., $\frac{\psi_{m-2}(\mathbf{x}_{0},\bm{\Pi}_{0})}{a_{m-1,0}}\geq 0),a_{m-1,t}>0,$ then we have $\psi_{m-2}(\mathbf{x}_{t},\bm{\Pi}_{t})\geq 0$ (This proof is similar to the proof in Rem. 1), and also $\frac{\psi_{m-2}(\mathbf{x}_{t},\bm{\Pi}_{t})}{a_{m-1,t}}\geq 0.$ Based on (22), since $(\mathbf{x}_{0},\bm{\Pi}_{0})\in\mathcal{C}_{m-3},a_{m-2,t}>0$ then similarly we have $\psi_{m-3}(\mathbf{x}_{t},\bm{\Pi}_{t})\geq 0$ and $\frac{\psi_{m-3}(\mathbf{x}_{t},\bm{\Pi}_{t})}{a_{m-2,t}}\geq 0,\forall t\in\mathbb{N}_{\geq 0}$. Repeatedly, we have $\psi_{0}(\mathbf{x}_{t},\bm{\Pi}_{t})\geq 0$ and $\frac{\psi_{0}(\mathbf{x}_{t},\bm{\Pi}_{t})}{a_{1,t}}\geq 0,\forall t\in\mathbb{N}_{\geq 0}$. Therefore the intersection of sets $\mathcal{C}_{0},\dots,\mathcal{C}_{m-1}$ is forward invariant and $h(\mathbf{x}_{t})=\frac{\psi_{0}(\mathbf{x}_{t},\bm{\Pi}_{t})}{a_{1,t}}\geq 0$ is ensured for all $t\in\mathbb{N}_{\geq 0}$. ∎
@@ IV-B Risk Sensitive Forward Invariance
In the presence of stochastic uncertainty $\mathbf{w}_{t}$, ensuring absolute invariance or safety is probably not feasible. Moreover, the uncertainties propagate dynamically over multiple steps until any component of the control input $\mathbf{u}_{t}$ explicitly appears, as shown in Def. 1 and Eq. (13). In this work, we propose ensuring the forward invariance of sets defined by DAVCBFs in the dynamic coherent risk measure sense to guarantee stochastic safety.
@@
Since $\rho_{0:t}$ generally represents a method for measuring risk, equation (26) implies risk-measured forward invariance. Moreover, $\rho_{0:t}$ can be expressed in a risk-sensitive form, such as CVaR, as shown in (12). Therefore, equation (26) can also imply risk-sensitive forward invariance.
@@ IV-D Optimal Control with RACBFs
In (30), $0\leq t\leq t_{T},t\in\mathbb{N}$, $\mathbf{x}_{e}$ denotes a reference state, $\mathbf{a}_{w}$ is the vector to which we hope auxiliary input $\bm{\nu}_{t}$ converges, and $Q,R,W$ are weight matrices with positive diagonal entries. $P$ is a positive weight scalar for CVaR and the last term of (30) is a transformation of Eq. (12). To enforce RACBFs in (27) with CVaR, we construct reformulated RACBF constraints as follows:
@@
We formulate control bounds $\mathbf{u}_{t}\in\mathcal{U}\subset\mathbb{R}^{q}$, along with the DHOCBFs and reformulated RACBFs introduced in Eq. (24) and Eq. (31) as constraints in the nonlinear optimization problem, minimizing the cost function given in Eq. (30).
@@
Another approach to incorporating auxiliary variables into DCBFs is by multiplying a positive $a_{i,t}$ in front of each class $\kappa$ function. Using this method, we rewrite Eq. (32) as:
@@
Compared to Eq. (32), we observe that $(1-\lambda_{i}a_{i,t})$ has a range of $[-\infty,1]$, while $\frac{a_{i,t}}{a_{i,t+1}}(1-\lambda_{i})\phi_{i-1}$ ranges from $[0,+\infty]$. The form of Eq. (32) maximizes the flexibility in adjusting the parameters of the class $\kappa$ function while eliminating negative parameters that may compromise safety.

### v2

@@ Abstract
This paper addresses the challenge of ensuring safety in stochastic control systems with high-relative-degree constraints, while maintaining feasibility and mitigating conservatism in risk evaluation. Control Barrier Functions (CBFs) provide an effective framework for enforcing safety constraints in nonlinear systems. However, existing methods struggle with feasibility issues and multi-step uncertainties. To address these challenges, we introduce Risk-aware Adaptive CBFs (RACBFs), which integrate Discrete-time Auxiliary-Variable adaptive CBFs (DAVCBFs) with coherent risk measures. DAVCBFs introduce auxiliary variables to improve the feasibility of the optimal control problem, while RACBFs incorporate risk-aware formulations to balance safety and risk evaluation performance. By extending discrete-time high-order CBF constraints over multiple steps, RACBFs effectively handle multi-step uncertainties that propagate through the system dynamics. We demonstrate the effectiveness of our approach on a stochastic unicycle system, showing that RACBFs maintain safety and feasibility while reducing unnecessary conservatism compared to standard robust formulations of discrete-time CBF methods.
@@ II-A Discrete-Time High-Order CBFs
Let the output mapping be $h:\mathbb{R}^{n}\rightarrow\mathbb{R}$. The output $y_{t}=h(\mathbf{x}_{t})$ of system (1) is said to have relative degree $m$ if
@@
i.e., $m$ is the number of steps (delay) in the output $y_{t}$ in order for any component of the control input $\mathbf{u}_{t}$ to explicitly appear ($\textbf{0}_{q}$ is the zero vector of dimension $q$).
@@
where $\alpha_{i}(\cdot)$ is a linear function and $0<\gamma_{i}\leq 1,i\in\{1,\dots,m\}$.
@@ II-B Coherent Risk Measure
The one-step conditional risk measures $\rho_{t}:\mathcal{H}_{t+1}\to\mathcal{H}_{t}$, $t=1,\dots,N-1$, as in (9), are called coherent risk measures if they satisfy the following properties:
@@
Conditional Value-at-Risk (CVaR) Let $h\in\mathcal{H}$ be a stochastic variable where higher values indicate safer performance. Since we aim to prevent $h$ from becoming negative (as seen in Eqn. (4)), we treat negative values of $h$ as a loss and measure left-tail risk using Conditional Value-at-Risk (CVaR) or Value-at-Risk (VaR) at the lower quantile. For a confidence level $\beta\in(0,1)$, VaR is defined as:
@@ III Problem Formulation and Approach
A control policy is feasible if all constraints derived from safety requirement and control limitation requirement are satisfied and mutually non-conflicting for all $0\leq t\leq t_{T}$. In this paper, we consider the following problem:
@@
Find a feasible control policy for system (13) and cost (14) is minimized.
@@ IV-A Enhancing Optimization Feasibility with DAVCBFs
where $\epsilon$ is a positive constant which can be infinitely small, the auxiliary input $\nu_{i,t}$ will appear in $\varphi_{i,0}(\bm{\pi}_{i,t+j_{m}})$ and $j_{m}=m+1-i$.
@@
Eqn. (19) and $\varphi_{i,m+1-i}(\bm{\pi}_{i,t})\geq\epsilon$ are equivalent. If this inequality constraint is satisfied, we have $\bigtriangleup\varphi_{i,m-i}(\bm{\pi}_{i,t})+\alpha_{i,m+1-i}(\varphi_{i,m-i}(\bm{\pi}_{i,t}))\geq\epsilon$. This can be rewritten as
@@
$\alpha_{j}(\psi_{j-1}(\mathbf{x}_{t},\bm{\Pi}_{t}))\leq\psi_{j-1}(\mathbf{x}_{t},\bm{\Pi}_{t})$, $\forall(\mathbf{x}_{t},\bm{\Pi}_{t})\in\mathcal{C}_{0}\cap,...,\cap\mathcal{C}_{m-1}$ and each $a_{i,t}>0,i\in\{1,\dots,m\}$. Note that the input $\mathbf{u}_{t}$ and auxiliary input $\bm{\nu}_{t}$ will appear in $\psi_{0}(\mathbf{x}_{t+m},\bm{\Pi}_{t+m})$ and Eqn. (25) is equivalent to $\psi_{m}(\mathbf{x}_{t},\bm{\Pi}_{t})\geq 0$.
@@
Given a DAVCBF $h(\mathbf{x})$ from Def. 5 with corresponding sets $\mathcal{C}_{0},\dots,\mathcal{C}_{m-1}$ defined by (23), if $(\mathbf{x}_{0},\bm{\Pi}_{0})\in\mathcal{C}_{0}\cap\dots\cap\mathcal{C}_{m-1},$ then if there exists a Lipschitz continuous controller $(\mathbf{u}_{t},\bm{\nu}_{t})$ that satisfies the constraint in (25) and also ensures $(\mathbf{x}_{t},\bm{\Pi}_{t})\in\mathcal{C}_{m-1}$ for all $t\in\mathbb{N}_{\geq 0}$, then $\mathcal{C}_{0}\cap\dots\cap\mathcal{C}_{m-1}$ will be rendered forward invariant for system (1), $i.e.,(\mathbf{x}_{t},\bm{\Pi}_{t})\in\mathcal{C}_{0}\cap\dots\cap\mathcal{C}_{m-1},\forall t\in\mathbb{N}_{\geq 0}$. Moreover, $h(\mathbf{x}_{t})\geq 0$ is ensured for all $t\in\mathbb{N}_{\geq 0}$
@@
If $h(\mathbf{x})$ is a DAVCBF, then satisfying the constraint in (25) while ensuring $(\mathbf{x}_{t},\bm{\Pi}_{t})\in\mathcal{C}_{m-1}$ for all $t\in\mathbb{N}_{\geq 0}$ is equivalent to making $\psi_{m-1}(\mathbf{x}_{t},\bm{\Pi}_{t})\geq 0,\forall t\in\mathbb{N}_{\geq 0}$. Since $a_{m,t}>0$, we have $\frac{\psi_{m-1}(\mathbf{x}_{t},\bm{\Pi}_{t})}{a_{m,t}}\geq 0.$ Based on (22), since $(\mathbf{x}_{0},\bm{\Pi}_{0})\in\mathcal{C}_{m-2}$ (i.e., $\frac{\psi_{m-2}(\mathbf{x}_{0},\bm{\Pi}_{0})}{a_{m-1,0}}\geq 0),a_{m-1,t}>0,$ then we have $\psi_{m-2}(\mathbf{x}_{t},\bm{\Pi}_{t})\geq 0$ (This proof is similar to the proof in Rem. 1), and also $\frac{\psi_{m-2}(\mathbf{x}_{t},\bm{\Pi}_{t})}{a_{m-1,t}}\geq 0.$ Based on (22), since $(\mathbf{x}_{0},\bm{\Pi}_{0})\in\mathcal{C}_{m-3},a_{m-2,t}>0$ then similarly we have $\psi_{m-3}(\mathbf{x}_{t},\bm{\Pi}_{t})\geq 0$ and $\frac{\psi_{m-3}(\mathbf{x}_{t},\bm{\Pi}_{t})}{a_{m-2,t}}\geq 0,\forall t\in\mathbb{N}_{\geq 0}$. Repeatedly, we have $\psi_{0}(\mathbf{x}_{t},\bm{\Pi}_{t})\geq 0$ and $\frac{\psi_{0}(\mathbf{x}_{t},\bm{\Pi}_{t})}{a_{1,t}}\geq 0,\forall t\in\mathbb{N}_{\geq 0}$. Therefore the intersection of sets $\mathcal{C}_{0},\dots,\mathcal{C}_{m-1}$ is forward invariant and $h(\mathbf{x}_{t})=\frac{\psi_{0}(\mathbf{x}_{t},\bm{\Pi}_{t})}{a_{1,t}}\geq 0$ is ensured for all $t\in\mathbb{N}_{\geq 0}$. ∎
@@ IV-B Risk Sensitive Forward Invariance
In the presence of stochastic uncertainty $\mathbf{w}_{t}$, ensuring absolute invariance or safety is probably not feasible. Moreover, the uncertainties propagate dynamically over multiple steps until any component of the control input $\mathbf{u}_{t}$ explicitly appears, as shown in Def. 1 and Eqn. (13). In this work, we propose ensuring the forward invariance of sets defined by DAVCBFs in the dynamic coherent risk measure sense to guarantee stochastic safety.
@@
Since $\rho_{0:t}$ generally represents a method for measuring risk, Eqn. (26) implies risk-measured forward invariance. Moreover, $\rho_{0:t}$ can be expressed in a risk-sensitive form, such as CVaR, as shown in (12). Therefore, Eqn. (26) can also imply risk-sensitive forward invariance.
@@ IV-D Optimal Control with RACBFs
In (30), $0\leq t\leq t_{T},t\in\mathbb{N}$, $\mathbf{x}_{e}$ denotes a reference state, $\mathbf{a}_{w}$ is the vector to which we hope auxiliary input $\bm{\nu}_{t}$ converges, and $Q,R,W$ are weight matrices with positive diagonal entries. $P$ is a positive weight scalar for CVaR and the last term of (30) is a transformation of Eqn. (12). To enforce RACBFs in (27) with CVaR, we construct reformulated RACBF constraints as follows:
@@
We formulate control bounds $\mathbf{u}_{t}\in\mathcal{U}\subset\mathbb{R}^{q}$, along with the DHOCBFs and reformulated RACBFs introduced in Eqn. (24) and Eqn. (31) as constraints in the nonlinear optimization problem, minimizing the cost function given in Eqn. (30).
@@
Another approach to incorporating auxiliary variables into DCBFs is by multiplying a positive $a_{i,t}$ in front of each class $\kappa$ function. Using this method, we rewrite Eqn. (32) as:
@@
Compared to Eqn. (32), we observe that $(1-\lambda_{i}a_{i,t})$ has a range of $[-\infty,1]$, while $\frac{a_{i,t}}{a_{i,t+1}}(1-\lambda_{i})\phi_{i-1}$ ranges from $[0,+\infty]$. The form of Eqn. (32) maximizes the flexibility in adjusting the parameters of the class $\kappa$ function while eliminating negative parameters that may compromise safety.